OpenStudy (anonymous):

Simplify a - {b - [c - (d - e) - f] - g} A) a + b + c - d + e - f + g B) a - b + c - d + e - f + g C) a + b + c + d + e - f + g D) a - b + c - d + e + f - g

OpenStudy (anonymous):

dis wingspan

OpenStudy (animal_lover36):

Use the distributive property to remove parentheses, then brackets, and then finally braces: a - {b - [c - (d - e) - f] - g} = a - {b - [c - d + e - f] - g} = a - {b - c + d - e + f - g} = a - b + c - d + e - f + g.

OpenStudy (anonymous):

so its A @animal_lover36

OpenStudy (animal_lover36):

I believe so.